OpenStudy (anonymous):

Solve the following quadratic equation using the square root property: (y + 6)^2 = -4 Only fill in one of the tables.

OpenStudy (anonymous):

i give medals

OpenStudy (anonymous):

@Venny

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I don't know what "square root property" is but the issue is that if you square root both sides from the beginning you'll get an imaginary number since the square root of -4 is equal to 2i. The first thing I'd do is distribute (y+6)^2 which turns out to equal (y^2) + 12y + 36 = -4. But I don't know if this is not doing it using the "square root property" way like you are asking.

OpenStudy (anonymous):

yea this pellet sucks

OpenStudy (anonymous):

pellet lol

OpenStudy (anonymous):

ok i guess i can cuss on here sh*it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Hmm I think the answer actually does involve imaginary numbers. When I use my calculator to solve the polynomial y^2 + 12y + 40 it gives me answers, both with imaginary numbers. I get -6-2i and -6+2i

jimthompson5910 (jim_thompson5910):

The square root property is the idea where you take the square root of both sides (to undo squaring) \[\Large (y + 6)^2 = -4\] \[\Large \sqrt{(y + 6)^2} = \pm \sqrt{-4}\] \[\Large y+6 = \pm \sqrt{-4}\] \[\Large y+6 = \pm 2i\] \[\Large y = 6 \pm 2i\] \[\Large y = 6 + 2i \ \text{ or } y = 6 - 2i\]

jimthompson5910 (jim_thompson5910):

Where \[\Large i = \sqrt{-1}\]

OpenStudy (anonymous):

it wont let me put I in because I can only type in so much

OpenStudy (anonymous):

You can't put in -6+2i or -6-2i?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you can't put in i, then just leave it as \[\sqrt{-4}\]

OpenStudy (anonymous):

So you should get \[-6-\sqrt{-4}\] and \[-6+\sqrt{-4}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

What's the second thing? I can't read your hand writing. The first one is \[\sqrt{-4}\] but what is the second one?

OpenStudy (anonymous):